OpenStudy (anonymous):

8=4(r+4)? if r=1 how do you check?

OpenStudy (anonymous):

replace where ever you see r with 1 and do the math calculation does 8 = 4(1+4)?

OpenStudy (anonymous):

8=4(1+4) 8=4*1+4*4 8=4+16 8≠20 (1 is not a solution)